OpenStudy (anonymous):

Find the inverse of the one-to-one function. f(x) = 2x+5/7 A. f-1(x) =7/2x+5 B. f-1(x) =7/2x-5 D. f-1(x) =7x-5/2

OpenStudy (anonymous):

you want the answer or how to find it ?

OpenStudy (anonymous):

We can't give answers, we are supposed to help the student with the method.

OpenStudy (anonymous):

solve for x and change to y

OpenStudy (anonymous):

f(x) = 2x+5/7 How can we solve that for x?

OpenStudy (anonymous):

I did it myself and got y=2x+5/2 but that was wrong But this is what I did y=2x+5/7 x=2y+5/7 y=7x+5/2 I don't know what I did wrong so no I want to know how to do it

OpenStudy (anonymous):

you dont change y for x in the beginning

OpenStudy (anonymous):

you have to solve for x first then change at the end

OpenStudy (anonymous):

for some reason i am really stuck do I make them fractions

OpenStudy (anonymous):

like this y= 2x/7 + 5/7

OpenStudy (anonymous):

or do I multiply the top by 7 like this y=14x + 35

OpenStudy (anonymous):

u subtract 5/7 first

OpenStudy (anonymous):

so 5/7 = 2x

OpenStudy (anonymous):

y-5/7

OpenStudy (anonymous):

so it is like this y-5/7 = 2x

OpenStudy (anonymous):

ya and you solve for x

OpenStudy (anonymous):

so then i get y - 5/14 = x

OpenStudy (anonymous):

y/2-5/14

OpenStudy (anonymous):

or 2y - 5/14 = x

OpenStudy (anonymous):